The number of mole of nitrogen in one litre of air containing 10% nitrogen by volume , under standard conditions is

Grade:11

1 Answers

Ashutosh Sharma
8 years ago
at STP 22.4 litre of nitrogen=1 mole
as the total volume of solution is= 1L
So for 10% nitrogen volume will be= 0.1L
as 22.4L=1mole
1L= 1/22.4 mole
therefore, in 0.1L= (1/22.4)x0.1=0.0044 mole.